OpenStudy (anonymous):

Describe how the graph of y= x^2 can be transformed to the graph of the given equation. y = x^2 - 12

OpenStudy (johnweldon1993):

 So as we can see...when we subtract a number from the function..we are moving the entire function down "number" units So here....to get from y = x^2 to y = x^2 - 12 We just need to move the graph of y = x^2 down 12 units

OpenStudy (anonymous):

oh okay thank you @johnweldon1993

OpenStudy (johnweldon1993):

No problem!